OpenStudy (anonymous):

Multiply and simplify radical 3 times the quantity 4+radical 12. A. the square root of 12+6 B. 4 and the square root of 3+6 C. 2 and the square root of 3+6 and the square root of 3

OpenStudy (anonymous):

\[\sqrt{3(4 +}\sqrt{12)} \] is the equation. The answer choices are: A. \[\sqrt{12 + 6}\] B. \[4\sqrt{3+6}\] C. \[2\sqrt{3+6}\sqrt{3}\]

OpenStudy (anonymous):

well let's start by first multiplying the 3 out

OpenStudy (anonymous):

How do I do that? :/ Ugh. Math is so confusing for me! Would I multiply it by the 4 or multiply it by itself?

OpenStudy (anonymous):

when i am finished with my question I will help u out

OpenStudy (anonymous):

Thank you very much :)

OpenStudy (anonymous):

ok I am back

OpenStudy (ash2326):

\[\sqrt{3(4+\sqrt{12}}\] first you open the bracket that can be done by multiplying 3*(4+sqrt12) that's 12+3*sqrt12

OpenStudy (anonymous):

Alright sweet :) What should I do for this problem?

OpenStudy (anonymous):

wait I am gonna figure it out

OpenStudy (ash2326):

now see it's \[\sqrt{12+\sqrt(12)}\] NOW 12=4*3 SO WE HAVE SQRT(12+SQRT(4*3)) 2 IS THE SQUARE ROOT OF 4 SO IT CAN BE TAKEN OUT OF THE RADICAL NOW WE HAVE \[\sqrt(12+2\sqrt(3))\]

OpenStudy (anonymous):

the answer is c

OpenStudy (anonymous):

Do u want an expalnation

OpenStudy (anonymous):

I won't bite

OpenStudy (ash2326):

meow i think the question is \[\sqrt(3)(4+\sqrt(12))\]

OpenStudy (anonymous):

But c makes sense

OpenStudy (anonymous):

with the way she originally wrote it

OpenStudy (anonymous):

No that's how it says. The way I wrote it.

OpenStudy (ash2326):

so we'll get \[4\sqrt(3)+\sqrt(3)\sqrt(12)\] next we'll multiply sqrt(3) and sqrt(12) we'll get sqrt(36) ie.6 \[4\sqrt(3)+\sqrt(36)\] we get \[4\sqrt(3)+6\]